A study found that 68% of the population owns a home in a random sample of 150 households 92 owned a home at the a=0.01 level is there enough evidence to reject the claim

1
Expert's answer
2022-05-05T04:22:43-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.68"

"H_a:p\not=0.68"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a two-tailed test is "z_c = 2.5758."

The rejection region for this two-tailed test is "R = \{z: |z| > 2.5758\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{\dfrac{92}{150}-0.68}{\sqrt{\dfrac{0.68(1-0.68)}{150}}}""\approx-1.75035"

Since it is observed that "|z| = 1.75035<2.5758=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-1.75035)=0.080058," and since "p=0.080058>0.01=\alpha," it is concluded that the null hypothesis is notrejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than "0.68," at the "\alpha = 0.01" significance level.